Fitzgerald Industries has a new project available that requires an initial investment of $4.7 million. The project will provide unlevered cash flows of $850,000 per year for the next 20 years. The company will finance the project with a debt-value ratio of .3. The company’s bonds have a YTM of 6.2 percent. The companies with operations comparable to this project have unlevered betas of 1.21, 1.14, 1.36, and 1.31. The risk-free rate is 3.6 percent and the market risk premium is 6.8 percent. The tax rate is 21 percent. What is the NPV of this project?
Answer #1
1] Average unlevered beta of the industry = (1.21+1.14+1.36+1.31)/4 = 1.255
2] Levered beta for the project = 1.255*(1+(1-21%)*0.3/0.7 = 1.68
3] After tax cost of debt = 6.2%*(1-21%) = 4.90%
Cost of equity per CAPM = 3.6%+1.68*6.8% = 15.02%
WACC = 4.90%*30%+15.02%*70% = 11.98%
4] NPV = 850000*(1.1198^20-1)/(0.1198*1.1198^20)-4700000 = $ 16,56,994.63
Answer #2

where does 70% comes from?

answered by: wulong